# Neural Network Reprogrammability: A Unified Theme on Model Reprogramming, Prompt Tuning, and Prompt Instruction

### v1

@@ Abstract
As large-scale pre-trained foundation models continue to expand in size and capability, efficiently adapting them to specific downstream tasks has become increasingly critical. Despite substantial research progress, existing adaptation approaches have evolved largely in isolation, without a clear understanding of their interrelationships. This survey introduces neural network reprogrammability as a unifying framework that bridges mainstream model adaptation techniques–model reprogramming, prompt tuning, and prompt instruction–previously fragmented research areas yet converges on a shared principle: repurposing a pre-trained model by manipulating information at the interfaces while keeping the model parameters frozen. These methods exploit neural networks’ sensitivity to information manipulation on different interfaces, be it through perturbing inputs, inserting tokens into intermediate layers, or providing task-specific examples in context, to redirect model behaviors towards desired outcomes. Building on the concept of reprogrammability, we present a taxonomy that categorizes such information manipulation-based adaptation approaches across four key dimensions: manipulation format (fixed or learnable), location (interfaces where manipulations occur), operator (how they are applied), and output alignment requirement (post-processing needed to align outputs with downstream tasks). Notably, this framework applies consistently across data modalities, independent of specific model architectures. Moreover, viewing established techniques like in-context learning and chain-of-thought prompting through this lens reveals both their theoretical connections and practical distinctions. We further analyze remaining technical challenges and ethical considerations, positioning neural network reprogrammability as a fundamental paradigm for efficient model adaptation. We lastly identify promising research directions emerging from this integrative viewpoint.
@@ 1 Introduction
A more efficient paradigm is gaining traction. Rather than modifying the internal parameters, a growing body of work now focuses on adapting pre-trained models for downstream tasks (i.e., target tasks) without extensive (or, in some cases, any) parameter updates. As Figure 1 (b) shows, these approaches strategically manipulate the target task-specific inputs or contextual information provided to a fixed pre-trained model, and then align the model output to the requirement of the target task’s output space. In this paper, we term such methods reprogrammability-centric adaptation (RCA) and formally define the concept of reprogrammability in Sec. 2.1. The shift of adaptation focus in RCA methods leads to a compelling benefit: model adaptation costs (in terms of trainable parameters) become largely independent of model size, correlating instead with the complexity of the task itself. Namely, input and output of the target data.
@@ 1.2 Purpose: Unifying Disparate Approaches
To address these issues, this survey introduces Neural Network Reprogrammability as a unifying conceptual framework to bring clarity and structure to the diverse landscape of adaptation methods. We define reprogrammability as the inherent capability of pre-trained models to be repurposed for new tasks without internal parameter updates. Instead, adaptation is achieved by strategically manipulating the inputs (and/or auxiliary) information provided to the model (Elsayed et al., 2019; Chen, 2024) and the output model produces, effectively leveraging neural networks’ inherent input-output sensitivity (Lowd and Meek, 2005) to elicit desired task-specific behaviors.
Figure 3: Neural Network Reprogrammability as a unifying framework to systematically understand overlapped but previously fragmented research areas.
Under this framework, we synthesize three major lines of work: model reprogramming (MR) (Elsayed et al., 2019; Chen, 2024), prompt tuning (PT) (Lester et al., 2021; Li and Liang, 2021), and in-context learning (Brown et al., 2020; Kirillov et al., 2023), which we refer to as prompt instruction (PI) throughout this paper to avoid confusion with PT. Although often developed in parallel across different communities, we argue that they all rely on a common underlying principle that has yet to be formally recognized and consolidated. We thus group them under the framework reprogrammability-centric adaptation (RCA). This perspective reveals the mathematical and conceptual connections between seemingly disparate methods, showing them as variations on a common theme: the systematic exploitation of a neural network’s sensitivity to manipulations at its interfaces.
With this unified view (Figure 3), we aim to bridge previously isolated research areas and uncover overarching principles of adaptation that hold across different techniques, data modalities, and applications. By examining these methods through the common lens of reprogrammability, we aim to clarify their relationships, foster more precise scientific discourse, facilitate the discovery of fundamental mechanisms, and thereby accelerate the creation of more effective and accessible ways to adapt large-scale pre-trained models.
@@ 1.3 Scope and Contributions of this survey
Instead, this survey builds upon them by presenting a structured overview for adapting pre-trained models without updating their parameters. We introduce neural network reprogrammability as the core concept and a new perspective to unify what has become a fragmented and often terminologically confusing area of research. We connect parameter-efficient adaptation techniques used across diverse modalities, emphasizing their shared strategy of input/output manipulation rather than parameter modification. This survey makes four primary contributions:
  • 1)
    A Conceptual Unification of MR, PT, and PI. We show that model reprogramming, prompt tuning, and prompt instruction–techniques often developed independently–share fundamental principles. They all exploit pre-trained representations and guide model behaviors by carefully crafting inputs and outputs. We provide a conceptual framework that bridges these distinct lines of study.
  • 2)
  • 3)
  • 4)
    A Discussion of Open Challenges and Future Research Directions. We identify fundamental open questions regarding the theoretical foundations of reprogrammability, analyze limitations of current evaluation methods, and discuss emerging ethical issues, aiming to catalyze future research in this rapidly evolving field.
@@
While several recent surveys have covered related areas, such as prompting or parameter-efficient fine-tuning (Gu et al., 2023; Schulhoff et al., 2024; Zhang et al., 2023a; Chen, 2024), this survey distinguishes itself by offering a distinct perspective and presenting neural network reprogrammability as a central concept to unite different adaptation families (MR, PT, PI) that avoid parameter updates. Existing surveys typically focus on tracking advancements within a specific technique (e.g., prompting or model reprogramming) or application area. In contrast, we propose a more general framework (Section 2) that spans traditionally separate research communities, along with a systematic taxonomy (Section 3) that classifies RCA methods along four key dimensions. This helps reveal connections that remain potentially hidden when these approaches are studied in isolation (See Table 1 for a comparative summary).
@@ 2 A Unified View of Reprogrammability
Here, we define reprogrammability as the capacity of fixed neural network to be redirected toward tasks outside its original training task, by manipulations within its input and output spaces, such that the new task’s context can fit in the model’s existing learned feature space. This concept unifies a range of parameter-efficient adaptation techniques by highlighting their shared reliance on the model’s intrinsic sensitivity to input and/or output transformations. Later, in Sec. 3, we will show how existing model adaptation strategies fit this definition and can be categorized along four dimensions: (1) the format of the manipulation, (2) its location within the model, (3) the operator used, and (4) the need for output alignment.
@@ 2.1 From Vulnerability to Adaptation
Repurposing Sensitivity for Adaptation. This inherent input sensitivity, conventionally framed as a weakness, can however be reinterpreted as the enabler of neural network reprogrammability. The very mechanisms that allow adversarial examples to drastically alter model predictions–namely, the ability to find specific input directions that maximally change the output–can be repurposed. Thus, a seemingly detrimental vulnerability can be transformed beneficially. Instead of perturbing inputs to make a model fail (e.g., via gradient ascent on the loss function), we can engineer input transformations to guide a pre-trained model to perform tasks it was not originally designed or trained for. Such a transformation forms the foundation of neural network reprogrammability.
@@
This viewpoint marks a paradigm shift in model adaptation. While traditional transfer learning methods, e.g., fine-tuning (Pan and Yang, 2009) and domain adaptation (Long et al., 2016, 2018), focus on modifying model weights, with reprogrammability, models can be instead adapted by manipulating the data fed to a fixed function. This implies that the emphasis shifts from altering the model to transforming the data. Through careful design of transformations, we can effectively “reprogram” a model’s input-output mapping, without touching its core learned representations.
@@ 2.2 Formulation of Reprogrammability
  • Manipulation Operator $\tau:{\mathcal{X}}^{\rm T}\times\lambda\to{\mathcal{M}}$ that defines how manipulation is incorporated along with the target task data, and $\tau\in\{{\rm add},{\rm concat},{\rm param}\}$. Generally, the operator can be addition $\tau_{\rm add}({\mathbf{x}}^{\rm T},\lambda)={\mathbf{x}}^{\rm T}+\lambda$ in cases where ${\mathbf{x}}^{\rm T}$ and $\lambda$ are appropriately dimensioned, or concatenation $\tau_{\rm concat}({\mathbf{x}}^{\rm T},\lambda)=\left[{\mathbf{x}}^{\rm T};\lambda\right]$, denoting the concatenation along a specified dimension, or parametric transformation $\tau_{\rm param}({\mathbf{x}}^{\rm T},\lambda)=\lambda({\mathbf{x}}^{\rm T})$ with a function $\lambda(\cdot)$ that maps target input to a space potentially compatible with the input space of $f$.
@@
  • Identity mapping: when ${\mathcal{Y}}^{\rm S}$ is directly compatible with ${\mathcal{Y}}^{\rm T}$.
  • Structured extraction: $O_{\omega}({\mathbf{y}}^{\rm S})=\omega({\mathbf{y}}^{\rm S})$, where $\omega$ parses and/or isolates task-relevant patterns from the pre-trained model’s output.
  • Class remapping: $O_{\omega}({\mathbf{y}}^{\rm S})=\omega{\mathbf{y}}^{\rm S}$, where $\omega\in\{0,1\}^{|{\mathcal{Y}}^{\rm T}|\times|{\mathcal{Y}}^{\rm S}|}$ is a permutation matrix; in this case, $\omega$ is often non-parametric, determined by statistical relationships (Tsai et al., 2020; Chen et al., 2023a).
  • Linear projections: $O_{\omega}({\mathbf{y}}^{\rm S})=\omega{\mathbf{y}}^{\rm S}$, where $\omega\in{\mathbb{R}}^{|{\mathcal{Y}}^{\rm T}|\times|{\mathcal{Y}}^{\rm S}|}$; this can be both parametric (Tsao et al., 2024) or non-parametric (Cai et al., 2024a).
@@ 2.4 Reprogrammability as a Unified Paradigm
The reprogrammability framework extends naturally beyond standard visual and textual domains to emerging modalities. As a larger range of sensory modalities has been incorporated into modern AI systems, e.g., audio, tactile (Yen et al., 2023; Hung et al., 2023), the same principles continue to hold. Yang et al. (2021) studied how to reprogram acoustic models for time-series data, Melnyk et al. (2023) reprogrammed LLMs for protein infilling tasks, and Jing et al. (2023); Fang et al. (2023) adapted pre-trained graph neural networks for molecular property prediction. In addition, language models are shown to perform protein infilling tasks (Vinod et al., 2025) and time-series forecasting tasks (Jin et al., 2024). These examples demonstrate how the mathematical formulation of reprogrammability generalizes across fundamentally different data structures and representations.
@@ 3 A Taxonomy of Reprogrammability
Table 2: A comparative analysis of existing RCA literature through the lens of reprogrammability. Under interface ($\ell$), setup ${\mathcal{X}}^{\rm S}$, ${\mathcal{E}}$, and ${\mathcal{H}}$ denote manipulations that occur in input space, embedding space, and hidden space, respectively. Under configuration ($\lambda$), setup OP and FX means optimizable and fixed manipulations. Under operator ($\tau$), setup AD, CO, and PR refers to additive, concatenative, and parametric operators. Under mapping ($\omega$), setup SA, LE, RB, and ID represent statistics-based alignment, learning-based alignment, rule-based alignment, and identity mapping.
@@ 3.2 Manipulation Location: Where to Modify
Raw Input Space ${\mathcal{X}}^{\rm S}$ Manipulations: Universal Adaptability.
@@
Embedding Space ${\mathcal{E}}$ Manipulations: Balancing Access and Efficiency.
@@
Hidden Representation Space ${\mathcal{H}}$ Manipulations: Deep Modification.
@@ 3.4 Output Alignment Requirements
Rule-based alignment applies deterministic procedures to extract relevant information from model outputs, enabling structured prediction without additional training. Still, it depends on human-designed interpretation rules. MR rarely employs complex rule-based alignment, since the output spaces between source and target tasks are often significant. This strategy is central to many PI methods, with CoT (Wei et al., 2022b) and ScratchPad (Nye et al., 2021) applying structured parsing to extract reasoning paths from generated text. ReAct (Yao et al., 2023) and Reflexion (Shinn et al., 2023) interpreted structured outputs like action sequences with pattern-based extraction. Multimodal instruction models like GPT-4V (Achiam et al., 2023) define $O_{\omega}$ as a fixed function that implements parsing, extraction, or validation rules designed to interpret model outputs in task-specific contexts.
Statistical Alignment: Non-parametric Mapping.
@@
Learnable Alignment: Parametric Transformation.
Learnable alignment introduces parametric functions, typically neural networks, to transform source model outputs to target requirements, offering greater flexibility through the optimizable mapping matrix but requiring additional parameters and training data. In MR, this corresponds to replacing statistical alignment with a linear probe (Neekhara et al., 2022). In contrast, learnable alignment is more common in PT (Li and Liang, 2021; Gu et al., 2022), where a learnable classification head (e.g., verbalizer) was employed atop prompted representations.

### v2

@@ Abstract
As large-scale pre-trained foundation models continue to expand in size and capability, efficiently adapting them to specific downstream tasks has become increasingly critical. Despite substantial research progress, existing adaptation approaches have evolved largely in isolation, without a clear understanding of their interrelationships. This survey introduces neural network reprogrammability as a unifying framework that bridges mainstream model adaptation techniques–model reprogramming, prompt tuning, and prompt instruction–previously fragmented research areas yet converges on a shared principle: repurposing a pre-trained model by manipulating information at the interfaces while keeping the model parameters frozen. These methods exploit neural networks’ sensitivity to information manipulation on different interfaces, be it through perturbing inputs, inserting tokens into intermediate layers, or providing task-specific examples in context, to redirect model behaviors towards desired outcomes. Building on the concept of reprogrammability, we present a taxonomy that categorizes such information manipulation-based adaptation approaches across four key dimensions: manipulation format (fixed or learnable), location (interfaces where manipulations occur), operator (how they are applied), and output alignment requirement (post-processing needed to align outputs with downstream tasks). Notably, this framework applies consistently across data modalities, independent of specific model architectures. Moreover, viewing established techniques such as in-context learning and chain-of-thought prompting through this lens reveals both their theoretical connections and practical distinctions. We further analyze remaining technical challenges and ethical considerations, positioning neural network reprogrammability as a fundamental paradigm for efficient model adaptation. We lastly identify promising research directions emerging from this integrative viewpoint.
@@ 1 Introduction
A more efficient paradigm is gaining traction. Rather than modifying the internal parameters, a growing body of work now focuses on adapting pre-trained models for downstream tasks (i.e., target tasks) without extensive (or, in some cases, any) parameter updates. As Figure 1 (b) shows, these approaches strategically manipulate the target task-specific inputs or contextual information provided to a fixed pre-trained model, and then align the model output to the requirement of the target task’s output space. In this paper, we term such methods reprogrammability-centric adaptation (RCA) and formally define the concept of reprogrammability in Sec. 2.1. Importantly, the shift of adaptation focus in RCA methods leads to a compelling benefit: model adaptation costs (in terms of trainable parameters) become largely independent of model size, correlating instead with the complexity of the task itself. Namely, the input and output dimensionality of the target data.
@@ 1.2 Purpose: Unifying Disparate Approaches
To address these issues, this survey introduces Neural Network Reprogrammability as a unifying conceptual framework to bring clarity and structure to the diverse landscape of parameter-efficient adaptation methods. We define reprogrammability as the inherent capability of pre-trained models to be repurposed for new tasks without internal parameter updates; whilst the adaptation is achieved by strategically manipulating the inputs (and/or auxiliary) information provided to the model (Elsayed et al., 2019; Chen, 2024) and the output model produces, effectively leveraging neural networks’ inherent input-output sensitivity (Lowd and Meek, 2005) to elicit desired task-specific behaviors.
Figure 3: We introduce Neural Network Reprogrammability as a unifying framework to bring coherence to a set of model adaptation techniques that have often been studied in isolation. Importantly, this shared underlying principle applies regardless of pre-trained model architectures and data modalities.
Reprogrammability provides a coherent lens through which to view three major families of work: model reprogramming (MR) (Elsayed et al., 2019; Chen, 2024), prompt tuning (PT) (Lester et al., 2021; Li and Liang, 2021), and in-context learning (Brown et al., 2020; Kirillov et al., 2023), which we refer to as prompt instruction (PI) throughout this paper to avoid confusion with PT. Although often developed in parallel across different communities, we argue that they all rely on a common underlying principle that has yet to be formally recognized and consolidated. This commonality motivates grouping them under the unifying concept of reprogrammability-centric adaptation (RCA). This perspective reveals the connections between seemingly disparate methods, showing them as variations on a common theme: the systematic exploitation of a neural network’s sensitivity to manipulations at its interfaces.
With this unified view (Figure 3), we aim to bridge previously disparate research efforts by uncovering the overarching principles of adaptation that hold across different techniques, data modalities, and applications. By examining these methods through the common lens of neural network reprogrammability, we intend to clarify their relationships, foster more precise cross-community terminology and discussions, and stimulate deeper investigations into the core mechanisms. Ultimately, this should accelerate the creation of more effective, efficient, and broadly accessible means for adapting large-scale pre-trained models.
@@ 1.3 Scope and Contributions of this survey
While several recent surveys have covered related areas, such as prompting or parameter-efficient fine-tuning (Gu et al., 2023; Schulhoff et al., 2024; Zhang et al., 2023a; Chen, 2024), this survey distinguishes itself by offering a distinct perspective and presenting neural network reprogrammability as a central concept to unite different adaptation families (MR, PT, PI) that have no internal parameter updates. Existing surveys typically focus on tracking advancements within a specific technique (e.g., prompting or model reprogramming) or application area. In contrast, we propose a more general framework (Section 2) that spans traditionally separate research communities, along with a systematic taxonomy (Section 3) that classifies RCA methods along four key dimensions. This helps reveal connections that remain potentially hidden when these approaches are studied in isolation (See Table 1 for a comparative summary).
Contributions.
This paper presents a structured overview for adapting pre-trained models without updating their parameters. We introduce neural network reprogrammability as the core concept and a new perspective to unify what has become a fragmented and often terminologically confusing area of research. We connect parameter-efficient adaptation techniques used across diverse modalities, emphasizing their shared strategy of input/output manipulation rather than parameter modification. This survey makes four primary contributions:
  • 1)
    A Conceptual Unification. We show that model reprogramming, prompt tuning, and prompt instruction–techniques often developed independently–share fundamental principles. They all exploit pre-trained representations and guide model behaviors by carefully crafting inputs and outputs. We provide a conceptual framework that bridges these distinct lines of study.
  • 2)
  • 3)
  • 4)
    A Discussion of Open Challenges and Future Directions. We identify fundamental open questions regarding the theoretical foundations of reprogrammability, analyze limitations of current evaluation methods, and discuss emerging ethical issues, aiming to catalyze future research in this rapidly evolving field.
@@ 2 A Unified View of Reprogrammability
Here, we define reprogrammability as the capacity of fixed neural network to be redirected toward tasks outside its original training task, by manipulations within its input and output spaces, such that the new task’s context can fit in the pre-trained model’s learned feature space. This concept unifies a range of parameter-efficient adaptation techniques by highlighting their shared reliance on the model’s intrinsic sensitivity to input and/or output transformations. Later, in Sec. 3, we will show how existing model adaptation strategies fit this definition and can be categorized along four dimensions: (1) the format of the manipulation, (2) its location within the model, (3) the operator used, and (4) the need for output alignment.
@@ 2.1 From Vulnerability to Adaptation
Repurposing Sensitivity for Adaptation. This inherent input sensitivity, conventionally framed as a weakness, can however be reinterpreted as the enabler of neural network reprogrammability. The very mechanisms that allow adversarial examples to drastically alter model predictions–namely, the ability to find specific input directions that maximally change the output–can be repurposed. Thus, a seemingly detrimental vulnerability can be transformed beneficially. Instead of perturbing inputs to make a model fail (e.g., via gradient ascent on the loss function), we can engineer input transformations to guide a pre-trained model to perform tasks it was not originally designed or trained for. Such transformations eventually form the foundation of neural network reprogrammability
@@
This viewpoint marks a paradigm shift in model adaptation. While traditional transfer learning methods, e.g., fine-tuning (Pan and Yang, 2009) and domain adaptation (Long et al., 2016, 2018), focus on modifying model weights, with reprogrammability, models can be instead adapted by manipulating the data fed to a fixed function. This implies that the emphasis shifts from updating the model to transforming the data. Through careful design of transformations, we can effectively “reprogram” a model’s input-output mapping, without touching its core learned representations.
@@ 2.2 Formulation of Reprogrammability
  • Manipulation Operator $\tau:{\mathcal{X}}^{\rm T}\times\lambda\to{\mathcal{M}}$ that defines how manipulation is incorporated along with the target task data, and $\tau\in\{{\rm add},{\rm concat},{\rm param}\}$. Generally, the operator can be addition $\tau_{\rm add}({\mathbf{x}}^{\rm T},\lambda)={\mathbf{x}}^{\rm T}+\lambda$ in cases where ${\mathbf{x}}^{\rm T}$ and $\lambda$ are appropriately dimensioned, or concatenation $\tau_{\rm concat}({\mathbf{x}}^{\rm T},\lambda)=\left[{\mathbf{x}}^{\rm T};\lambda\right]$, denoting the concatenation along a specified dimension, or parametric transformation $\tau_{\rm param}({\mathbf{x}}^{\rm T},\lambda)=\lambda({\mathbf{x}}^{\rm T})$ with a function $\lambda(\cdot)$ that maps target input to a space compatible with the input space of the corresponding interface.
@@
  • Identity Mapping. The simplest approach, used when ${\mathcal{Y}}^{\rm S}$ is directly compatible with ${\mathcal{Y}}^{\rm T}$, i.e., ${\mathcal{Y}}^{\rm S}={\mathcal{Y}}^{\rm T}$, requiring no explicit alignment transformation. This is common in generative tasks where the model directly produces outputs in the desired format.
  • Structured Alignment. This involves a mapping $O_{\omega}({\mathbf{y}}^{\rm S})=\omega({\mathbf{y}}^{\rm S})$, where $\omega(\cdot)$ is typically a pre-defined, fixed function (e.g., a rule-based parser or extractor), aiming to isolate or reformat task-relevant information from the raw output of the pre-trained model, often employed for tasks requiring specific output structures.
  • Statistical Alignment. This involves a mapping $O_{\omega}({\mathbf{y}}^{\rm S})=\omega{\mathbf{y}}^{\rm S}$, where $\omega\in\{0,1\}^{|{\mathcal{Y}}^{\rm T}|\times|{\mathcal{Y}}^{\rm S}|}$ is a binary matrix. The matrix $\omega$ is determined non-parametrically based on statistical correspondences between the model’s source predictions and the target labels, known as label mapping (Tsai et al., 2020; Chen et al., 2023a).
  • Linear Alignment. This also uses a matrix transformation $O_{\omega}({\mathbf{y}}^{\rm S})=\omega{\mathbf{y}}^{\rm S}$, but here $\omega\in{\mathbb{R}}^{|{\mathcal{Y}}^{\rm T}|\times|{\mathcal{Y}}^{\rm S}|}$ is a real-valued matrix. The matrix $\omega$ can be obtained either by gradient-based backpropagation (Tsao et al., 2024), e.g., optimizing a linear probe trained on top of the frozen pre-trained model’s outputs (Tsao et al., 2024); or statistically estimated without backpropagation, for example, by estimating the optimal probabilistic mapping matrix from observed statistics of source predictions and target labels (Cai et al., 2024a) [^2] We note that (Cai et al., 2024a) also relies on statistics and does not rely on gradient-based optimization. Still, here we distinguish it from the “statistical alignment” category as $\omega$ is not restricted to binary values and can represent more precise linear transformations..
@@ 2.3 Revisit Existing Manifestations of Reprogrammability
Figure 5: Examples of how reprogrammability manifests across different adaptation methodologies, data modalities, and downstream tasks. (a) MR that repurposes a pre-trained image classifier for a new image classification task. (b) PT that repurposes a pre-trained language generator for a new text generation task.
@@ 2.4 Reprogrammability as a Unified Paradigm
The reprogrammability framework extends naturally beyond standard visual and textual domains to emerging modalities. As a larger range of sensory modalities has been incorporated into modern AI systems, e.g., audio, tactile (Yen et al., 2023; Hung et al., 2023), the same principles continue to hold. Yang et al. (2021) studied how to reprogram acoustic models for time-series data, and Jing et al. (2023); Fang et al. (2023) adapted pre-trained graph neural networks for molecular property prediction. In addition, language models are shown to perform protein infilling tasks (Melnyk et al., 2023; Vinod et al., 2025) and time-series forecasting tasks (Jin et al., 2024). These examples illustrate how the underlying formulation of reprogrammability generalizes across fundamentally different data structures and representations.
@@ 3 A Taxonomy of Reprogrammability
Table 2: A comparative analysis of existing RCA literature through the lens of reprogrammability. Under interface ($\ell$), setup ${\mathcal{X}}^{\rm S}$, ${\mathcal{E}}$, and ${\mathcal{H}}$ denote manipulations that occur in input space, embedding space, and hidden space, respectively. The columns under configuration ($\lambda$): setup OP and FX means optimizable and fixed manipulations. The columns under operator ($\tau$): setup AD, CO, and PR refers to additive, concatenative, and parametric operators. The columns under mapping ($\omega$): setup SA, LA, RA, and ID represent statistical alignment, linear alignment, rule-based (i.e., structured) alignment, and identity mapping.
@@ 3.2 Manipulation Location: Where to Modify
Raw Input Space ${\mathcal{X}}^{\rm S}$ Manipulations.
@@
Embedding Space ${\mathcal{E}}$ Manipulations.
@@
Hidden Representation Space ${\mathcal{H}}$ Manipulations.
@@ 3.4 Output Alignment Requirements
Rule-based alignment applies deterministic procedures to extract relevant information from model outputs, enabling structured prediction without additional training. Still, it depends on human-designed interpretation rules. MR rarely employs complex rule-based alignment, since the output spaces between source and target tasks are often significant. This strategy is central to many PI methods, with CoT (Wei et al., 2022b) and ScratchPad (Nye et al., 2021) applying structured parsing to extract reasoning paths from generated text. ReAct (Yao et al., 2023) and Reflexion (Shinn et al., 2023) interpreted structured outputs like action sequences with pattern-based extraction. Multimodal instruction models like GPT-4V (Achiam et al., 2023) define $O_{\omega}$ as a pre-defined function that implements parsing, extraction, or validation rules designed to interpret model outputs in task-specific contexts.
Statistical Alignment: Non-parametric Label Mapping.
@@
Linear Alignment: Parametric Transformation.
Linear alignment introduces parametric functions, typically neural networks, to transform source model outputs to target requirements, offering greater flexibility through the optimizable mapping matrix but requiring additional parameters and training data. In MR, this corresponds to replacing statistical alignment with a linear probe (Neekhara et al., 2022). In contrast, learnable alignment is more common in PT (Li and Liang, 2021; Gu et al., 2022), where a learnable classification head (e.g., verbalizer) was employed atop prompted representations.